# Conformal Prediction in Hierarchical Classification

### v1

@@ Abstract
Conformal prediction has emerged as a widely used framework for constructing valid prediction sets in classification and regression tasks. In this work, we extend the split conformal prediction framework to hierarchical classification, where prediction sets are commonly restricted to internal nodes of a predefined hierarchy, and propose two computationally efficient inference algorithms. The first algorithm returns internal nodes as prediction sets, while the second relaxes this restriction, using the notion of representation complexity, yielding a more general and combinatorial inference problem, but smaller set sizes. Empirical evaluations on several benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving nominal coverage.
Machine Learning, ICML
@@
1 Introduction
In multi-class classification, a classifier can be uncertain about the predicted class label for a given test instance. In such cases, it can be beneficial to return set-valued predictions, i.e. sets of classes rather than individual classes. This is particularly relevant in hierarchical classification, where the class space is organized in a hierarchical structure, such as in medical diagnosis, where diseases are organized in a tree structure based on the International Classification of Diseases (ICD) (World Health Organization, 1978). In hierarchical classification, set-valued predictions are often limited to specific internal nodes in the hierarchy. Such sets have a clear semantic interpretation and can be constructed using efficient inference algorithms (Alex Freitas, 2007; Bi & Kwok, 2015; Rangwala & Naik, 2017; Yang et al., 2017; Wang et al., 2021; Valmadre, 2022). However, restricting predictions in this way can affect performance. For this reason, some authors allow any subset of classes as a prediction in hierarchical classification (Oh, 2017). Recently, a set-valued prediction framework for hierarchical classification that makes a compromise between the two aforementioned extremes has been proposed (Mortier et al., 2022). Compared to approaches that predict a single node of the hierarchy, this framework allows the user to restrict the representation complexity of a predicted set. The main idea is to return a restricted number of internal nodes of the hierarchy as candidate sets instead of a single node. However, this framework lacks a formal statistical guarantee and relies on well-calibrated probabilistic models.
Contributions. In this work, we build further upon the concept of representation complexity and extend the split conformal prediction framework to the hierarchical multi-class classification setting. Conformal prediction is a framework that provides valid and efficient set-valued predictions for any learning algorithm, without making any assumptions on the underlying data distribution (Vovk et al., 2005). In particular, given a trained (hierarchical) classifier, a desired coverage level $1-\alpha\in(0,1)$, $N$ calibration samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ and a test sample $(\bm{x}_{N+1},y_{N+1})$, drawn i.i.d. from an unknown distribution $P$, we would like to construct a prediction set $\hat{Y}\in 2^{\mathcal{Y}}$ with representation complexity $r$ for the test instance $(\bm{x}_{N+1},y_{N+1})$, such that the following marginal validity guarantee holds:
$\mathbb{P}\left[y_{N+1}\in\hat{Y}(\bm{x}_{N+1})\right]\geq 1-\alpha\,,\quad R_{\mathcal{T}}(\hat{Y})\leq r\,,$ (1)
where $R_{\mathcal{T}}(\hat{Y})$ denotes the representation complexity of the set $\hat{Y}$, given some tree structure $\mathcal{T}$. The probability above is taken over all $N+1$ samples and we require that (1) holds for any fixed $\alpha,N,r$ and $P$.
After reviewing some essentials on hierarchical probabilistic classification and conformal prediction in Section 2, we propose in Section 3 two algorithms that construct valid prediction sets, in the sense of satisfying (1). The first algorithm is designed for the restrictive case $r=1$, while the second extends to $r>1$. Moreover, we show that our algorithms possess distribution-free finite sample guarantees. Finally, in Section 4, we illustrate the notion of representation complexity and evaluate the proposed algorithms in terms of coverage and efficiency on a range of benchmark datasets.
@@ 1 Introduction
Figure 1: An example of a tree structure $\mathcal{T}$ with class space $\mathcal{Y}=\{1,\ldots,8\}$ and nodes $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{15}\}$. The root $v_{1}$ represents the class space $\mathcal{Y}$ and leaves $\{v_{8},\ldots,v_{15}\}$ represent the individual classes. The numbers in the leaf nodes represent the class probabilities for some instance $\bm{x}$.
2 Background and related work
@@ 2.1 Hierarchical multi-class classification
In a standard multi-class classification setting one assumes that training and test data are i.i.d. according to an unknown distribution $P(\bm{x},y)$ on $\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ some instance space (e.g. images, documents, etc.) and $\mathcal{Y}=\{c_{1},\ldots,c_{K}\}$ a class space consisting of $K$ classes. Probabilistic multi-class classifiers estimate the conditional class probabilities $P(\cdot\,|\,\bm{x})$ over $\mathcal{Y}$, with properties $\forall c\in\mathcal{Y}:0\leq P(c\,|\,\bm{x})\leq 1\,,\sum_{c\in\mathcal{Y}}P(c\,|\,\bm{x})=1\,.$ This distribution can be estimated using a wide range of well-known probabilistic models, such as logistic regression, linear discriminant analysis, gradient boosting trees or neural networks with a softmax output layer. At prediction time, set-valued prediction algorithms, such as split conformal prediction, return sets $\hat{Y}$ that are subsets of $\mathcal{Y}$. The probability mass of such a set can be computed as $P(\hat{Y}\,|\,\bm{x})=\sum_{c\in\hat{Y}}P(c\,|\,\bm{x})$.
However, in this paper we will consider a hierarchical multi-class classification setting. Hence, we assume that a domain expert has defined a hierarchy over the class space, in the form of a tree structure $\mathcal{T}$ that contains in general $M$ nodes. $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{M}\}$ will denote the set of nodes and every node identifies a set of classes. As special cases, the root $v_{1}$ represents the class space $\mathcal{Y}$, and the leaves represent individual classes – see Figure 1 for a simple example. In hierarchical classification, one typically makes the strong restriction $\hat{Y}\in\mathcal{V}_{\mathcal{T}}$ for predicted sets – see e.g. Bi & Kwok (2015). The probability mass $P(v\,|\,\bm{x})$ of such a set can be computed using the chain rule of probability:
@@
where $\mathrm{Path}(v)$ is a set of nodes on the path connecting the node $v$ and the root of the tree structure. $\mathrm{pa}(v)$ gives the parent of node $v$ and $P(v\,|\,\mathrm{pa}(v),\bm{x})$ represents the branch probability of node $v$ given its parent $\mathrm{pa}(v)$. Note that for the root node $v_{1}$ one has $P(v_{1}\,|\,\mathrm{pa}(v_{1}),\bm{x})=1$. In Figure 1, the branch probabilities of the root node $v_{1}$ are given by $P(v_{2}\,|\,v_{1},\bm{x})=0.485$ and $P(v_{3}\,|\,v_{1},\bm{x})=0.515$. In order to estimate the branch probabilities, one can train any multi-class probabilistic classifier in each internal node of the tree. Classical models of that kind include nested dichotomies (Fox, 1997; Frank & Kramer, 2004; Melnikov & Hüllermeier, 2018), conditional probability estimation trees (Beygelzimer et al., 2009) and probabilistic classifier trees (Dembczyński et al., 2016). In neural networks with a hierarchical softmax output layer, all nodes are trained simultaneously (Morin & Bengio, 2005).
@@ 2.2 Inference in hierarchical classification
In this work, we do not focus on the training algorithms. Instead, we assume that a probabilistic model $\hat{P}$ has been estimated, either with classical models or using a hierarchical factorization as in (2), and we focus on the prediction task. In particular, we would like to predict sets $\hat{Y}\in 2^{\mathcal{Y}}$ that satisfy (1), with a restriction on the representation complexity of the predicted set (Mortier et al., 2022). The representation complexity is defined as the minimal number of nodes needed to represent the set $\hat{Y}$ in the tree structure $\mathcal{T}$. More formally, let $\mathcal{S}_{\mathcal{T}}(\hat{Y})$ denote the set of all disjoint combinations of tree nodes that represent the set $\hat{Y}$:
@@
$R_{\mathcal{T}}(\hat{Y})=\min_{\hat{V}\in\mathcal{S}_{\mathcal{T}}(\hat{Y})}|\hat{V}|\,,$ (3)
with $|\hat{V}|$ the cardinality of $\hat{V}$. For example, the set $\hat{Y}=\{1,2,4,7,8\}$ of the hierarchy shown in Figure 1 has representation complexity three, because three nodes are needed to represent this set: $v_{4}$, $v_{7}$ and $v_{11}$.
@@ 2.3 Randomized nested prediction sets with split conformal prediction
Now we will review a general procedure for valid set-valued predictions in flat classification (i.e. ignoring the hierarchy), following the work of Gupta et al. (2022); Romano et al. (2020); Angelopoulos et al. (2020). Compared to traditional conformal prediction—which starts from the notion of a nonconformity score—this procedure departs from a sequence of nested set-valued predictions, where the set size depends on a threshold $\tau$. Furthermore, an independent calibration set is used to tune the threshold $\tau$ such that $(1-\alpha)$-coverage is guaranteed on future test samples. The use of a single calibration set is better known as split conformal prediction and gives rise to computationally efficient valid set-valued predictions (Papadopoulos et al., 2002; Lei et al., 2018).
More formally, let $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ be an i.i.d. sequence of $N$ samples from the unknown distribution $P$, and assume that these samples were not used for model training. Let $\hat{Y}(\bm{x},u,\tau):\mathcal{X}\times[0,1]\times\mathbb{R}\rightarrow 2^{\mathcal{Y}}$ be a set-valued predictor. In the spirit of Romano et al. (2020); Angelopoulos et al. (2020), the second argument $u$ represents a random draw from a uniform distribution $\mathcal{U}(0,1)$ and is included to allow for randomized prediction sets. The third argument $\tau$ is a threshold that controls the size of the predicted set. Assume that the sets are indexed by $\tau$ and are nested:
@@
$\tau^{*}=\inf\big{\{}\tau\in[0,1]:|\{i:y_{i}\in\hat{Y}(\bm{x}_{i},u_{i},\tau)\}|$
$\geq\lceil(1-\alpha)(N+1)\rceil\big{\}}\,.$ (5)
The set-valued predictor $\hat{Y}(\bm{x}_{N+1},u_{N+1},\tau^{*})$, with $\tau^{*}$ in (2.3), has marginal validity guarantees, as shown by the theorem below.
Theorem 2.1 (Marginal validity of nested conformal prediction (Angelopoulos et al., 2020)).
Assume an exchangeable sequence $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N+1}$ and let $\hat{Y}(\bm{x},u,\tau)$ be a set-valued predictor that satisfies (4). Furthermore, assume that $\exists\tau\in\mathbb{R}:\hat{Y}(\bm{x},u,\tau)=\mathcal{Y}$. Then, for $\tau^{*}$ in (2.3) and any $\alpha\in(0,1)$, the following marginal coverage guarantee holds:
@@
Moreover, if the sets grow smoothly in $\tau$, then the following upper bound holds:
$\mathbb{P}\big{[}y_{N+1}\in\hat{Y}(\bm{x}_{n+1},u_{N+1},\tau^{*})\big{]}\leq 1-\alpha+\frac{1}{N+1}\,.$
In the literature, several conformal prediction methods can be found that satisfy (4). For example, in Romano et al. (2020), the following set-valued predictor, called adaptive prediction sets (APS), is proposed:
@@
with $\hat{\rho}(y;\bm{x})=\sum_{y^{\prime}\in\mathcal{Y}}\hat{P}(y^{\prime}\,|\,\bm{x})\mathbbm{1}_{\hat{P}(y^{\prime}\,|\,\bm{x})>\hat{P}(y\,|\,\bm{x})}$ the probability mass of the labels more likely than $y$. At the heart of their method is the randomization term $u\cdot\hat{P}(y\,|\,\bm{x})$, which has been introduced in order to achieve exact nominal coverage. In addition, by means of the cumulative distribution $\hat{\rho}(y;\bm{x})$, this method allows to adapt effectively to complex data distributions, achieving a better conditional coverage compared to alternative approaches that rely on the mode of the distribution, such as the least ambiguous classifier proposed by Sadinle et al. (2019). In Angelopoulos et al. (2020), a variant of the above method has been proposed, called the regularized adaptive prediction sets (RAPS) method. By introducing a regularization term in (6), which needs to be tuned by means of additional tuning data, small set sizes are encouraged, in particular, when faced with noisy probability estimates for classes with low probability.
Alternative conformal prediction methods have been introduced, focusing on minimizing set size (Sadinle et al., 2019) and achieving approximate coverage across the entire feature space (Foygel Barber et al., 2021; Cauchois et al., 2021; Romano et al., 2019). In Angelopoulos et al. (2023), a conformal prediction method is proposed that controls the risk with respect to any bounded non-increasing loss function. Interestingly, this method is also applicable to the setting of hierarchical classification, i.e. through the use of the tree-distance loss, introduced in Bi & Kwok (2015) as a way of evaluating set-valued predictions in multi-label classification. However, set-valued predictions of that kind lack meaningful and practical interpretation, compared to methods that rely on the traditional notion of coverage. Furthermore, this method is not directly applicable to our work, since it does not allow the incorporation of representation complexity, given the properties of the loss function above.

### v3

@@ Abstract
Conformal prediction has emerged as a widely used framework for constructing valid prediction sets in classification and regression tasks. In this work, we extend the split conformal prediction framework to hierarchical classification, where prediction sets are commonly restricted to internal nodes of a predefined hierarchy, and propose two computationally efficient inference algorithms. The first algorithm returns internal nodes as prediction sets, while the second one relaxes this restriction. Using the notion of representation complexity, the latter yields smaller set sizes at the cost of a more general and combinatorial inference problem. Empirical evaluations on several benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving nominal coverage.
1 Introduction
In multi-class classification, a classifier can be uncertain about the predicted class label for a given test instance. In such cases, it can be beneficial to return set-valued predictions, i.e. sets of classes rather than individual classes. This is particularly relevant in hierarchical classification, where the class space is organised in a hierarchical structure, such as in medical diagnosis, where diseases are organised in a tree structure based on the International Classification of Diseases (ICD) (World Health Organization,, 1978).
In hierarchical classification, set-valued predictions are often restricted to internal nodes of the hierarchy. Such predictions have a clear semantic interpretation and can be computed using efficient inference algorithms (Alex Freitas,, 2007; Bi and Kwok,, 2015; Rangwala and Naik,, 2017; Yang et al.,, 2017; Wang et al.,, 2021; Valmadre,, 2022). However, a restriction of this kind can affect efficiency when the classifier is uncertain between classes in different branches of the hierarchy. In these cases, predicting a single internal node typically yields large and uninformative sets.
To address this, some approaches allow any subset of classes as predictions, which improves flexibility but comes at the cost of higher semantic complexity and reduced interpretability (Oh,, 2017). More recently, a set-based utility maximisation framework has been proposed that makes a compromise between the two aforementioned extremes by introducing the notion of representation complexity (Mortier et al.,, 2022). The assumption behind this notion is that inner nodes of a hierarchy have semantic meaning and can therefore be used as meaningful stand-alone predictions. By limiting the number of internal nodes used to represent a prediction, the proposed framework allows users to control the trade-off between efficiency and interpretability.
Figure 1: A sample image of the Lotus corniculatus species from the PlantCLEF 2015 dataset (Göeau et al.,, 2015).
To illustrate the usefulness of representation complexity, consider the PlantCLEF 2015 dataset (Göeau et al.,, 2015). This dataset consists of over one hundred thousand images representing 1,000 species of trees, herbs, and ferns native to the Western European region and is characterised by significant class ambiguity, making accurate predictions on species level often impossible. In Figure 1, an example image is shown corresponding to the Lotus corniculatus species. Using our proposed algorithms, two set-valued predictions were computed, with a restricted representation complexity of $r=1$ and $r\leq 3$, respectively. The former yields the root node of the hierarchy, i.e. all 1,000 species, and is completely uninformative. Increasing the representation complexity to three improves efficiency and yields the prediction $\{\emph{Lotus corniculatus},\emph{Tulipa sylvestris},\emph{Ficaria verna}\}$, i.e., three visually-related species including the true label.
Contributions. In this work, we build further upon the concept of representation complexity and extend the split conformal prediction framework to the hierarchical multi-class classification setting. Conformal prediction provides valid and efficient set-valued predictions for any learning algorithm, without making any assumptions on the underlying data distribution (Vovk et al.,, 2005). In particular, given a trained (hierarchical) classifier, a desired coverage level $1-\alpha\in(0,1)$, calibration samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$ and a test sample $(\bm{x}_{N+1},y_{N+1})$, drawn i.i.d. from an unknown distribution $P$, we would like to construct a prediction set $\hat{Y}\in 2^{\mathcal{Y}}$ with representation complexity $r$ for the test instance $(\bm{x}_{N+1},y_{N+1})$, such that the following marginal validity guarantee holds:
$\mathbb{P}\left[y_{N+1}\in\hat{Y}(\bm{x}_{N+1})\right]\geq 1-\alpha\,,\,\,\textrm{s.t.}\,\,R_{\mathcal{T}}(\hat{Y})\leq r\,$, (1)
@@ 1 Introduction
After reviewing some essentials on hierarchical probabilistic classification and conformal prediction in Section 2, we propose in Section 3 two algorithms that construct valid prediction sets, in the sense of satisfying (1). The first algorithm is designed for the restrictive case $r=1$, while the second extends to $r>1$. Moreover, we show that our algorithms possess distribution-free finite sample guarantees. Finally, in Section 4, we evaluate the proposed algorithms in terms of coverage and efficiency on a range of benchmark datasets.
@@
Figure 2: An example of a tree structure $\mathcal{T}$ with class space $\mathcal{Y}=\{1,\ldots,8\}$ and nodes $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{15}\}$. The root $v_{1}$ represents the class space $\mathcal{Y}$ and leaves $\{v_{8},\ldots,v_{15}\}$ represent the individual classes. The numbers in the leaf nodes represent the class probabilities for an instance $\bm{x}$.
2 Background, related work and formal description of existing concepts
@@ 2.1 Hierarchical multi-class classification
In a standard multi-class classification setting, one assumes that training and test data are i.i.d. according to an unknown distribution $P(\bm{x},y)$ on $\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ some instance space (e.g. images, documents, etc.) and $\mathcal{Y}=\{c_{1},\ldots,c_{K}\}$ a class space consisting of $K$ classes. Probabilistic multi-class classifiers estimate the conditional class probabilities $P(\cdot\,|\,\bm{x})$ over $\mathcal{Y}$, such that $0\leq P(c\,|\,\bm{x})\leq 1$ for all $c\in\mathcal{Y}$ and $\sum_{c\in\mathcal{Y}}P(c\,|\,\bm{x})=1\,.$ This distribution can be estimated using a wide range of well-known probabilistic models, such as logistic regression, linear discriminant analysis, gradient boosting trees or neural networks with a softmax output layer. At prediction time, set-valued prediction algorithms, such as split conformal prediction, return sets $\hat{Y}$ that are subsets of $\mathcal{Y}$. The probability mass of such a set can be computed as $P(\hat{Y}\,|\,\bm{x})=\sum_{c\in\hat{Y}}P(c\,|\,\bm{x})$.
In this paper, we will consider a hierarchical multi-class classification setting. Hence, we assume that a domain expert has defined a hierarchy over the class space, in the form of a tree structure $\mathcal{T}$ that in general contains $M$ nodes. $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{M}\}$ will denote the set of nodes and every node identifies a set of classes. As special cases, the root $v_{1}$ represents the class space $\mathcal{Y}$, and the leaves represent individual classes; see Figure 2 for a simple example. In hierarchical classification, the strong restriction $\hat{Y}\in\mathcal{V}_{\mathcal{T}}$ is typically made for predicted sets; see e.g. Bi and Kwok, (2015). The probability mass of such a set can be computed as $P(v\,|\,\bm{x})=\sum_{c\in v}P(c\,|\,\bm{x})$, or by using the chain rule of probability:
@@
where $\mathrm{Path}(v)$ is a set of nodes on the path connecting the node $v$ and the root of the tree structure. $\mathrm{pa}(v)$ gives the parent of node $v$ and $P(v\,|\,\mathrm{pa}(v),\bm{x})$ represents the branch probability of node $v$ given its parent $\mathrm{pa}(v)$. Note that for the root node $v_{1}$ one has $P(v_{1}\,|\,\mathrm{pa}(v_{1}),\bm{x})=1$. In Figure 2, the branch probabilities of the root node $v_{1}$ are given by $P(v_{2}\,|\,v_{1},\bm{x})=0.485$ and $P(v_{3}\,|\,v_{1},\bm{x})=0.515$. In order to estimate the branch probabilities, one can train any multi-class probabilistic classifier in each internal node of the tree. Classical models of that kind include nested dichotomies (Fox,, 1997; Frank and Kramer,, 2004; Melnikov and Hüllermeier,, 2018), conditional probability estimation trees (Beygelzimer et al.,, 2009) and probabilistic classifier trees (Dembczyński et al.,, 2016). In neural networks with a hierarchical softmax output layer, all nodes are trained simultaneously (Morin and Bengio,, 2005).
@@ 2.2 Inference in hierarchical classification
[…]
@@
[…]
@@ 2.3 Randomized nested prediction sets with split conformal prediction
Now we will review a general procedure for valid set-valued predictions in flat classification (i.e. ignoring the hierarchy), following the work of Gupta et al., (2022); Romano et al., (2020); Angelopoulos et al., (2020). Compared to traditional conformal prediction—which starts from the notion of a nonconformity score—this procedure departs from a sequence of nested set-valued predictions, where the set size depends on a threshold $\tau$. Furthermore, an independent calibration set is used to tune the threshold $\tau$ such that $(1-\alpha)$-coverage is guaranteed on future test samples. The use of a single calibration set is better known as split conformal prediction and gives rise to computationally efficient valid set-valued predictions (Papadopoulos et al.,, 2002; Lei et al.,, 2018).
More formally, let $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ be an i.i.d. sequence of $N$ samples from the unknown distribution $P$, and assume that these samples were not used for model training. Let $\hat{Y}(\bm{x},u,\tau):\mathcal{X}\times[0,1]\times\mathbb{R}\rightarrow 2^{\mathcal{Y}}$ be a set-valued predictor. In the spirit of Romano et al., (2020); Angelopoulos et al., (2020), the second argument $u$ represents a random draw from a uniform distribution $\mathcal{U}(0,1)$ and is included to allow for randomized prediction sets. The third argument $\tau$ is a threshold that controls the size of the predicted set. Assume that the sets are indexed by $\tau$ and are nested:
@@
$\tau^{*}=\inf\big\{\tau\in[0,1]:|\{i:y_{i}\in\hat{Y}(\bm{x}_{i},u_{i},\tau)\}|\\ \geq\lceil(1-\alpha)(N+1)\rceil\big\}\,.$ (5)
The set-valued predictor $\hat{Y}(\bm{x}_{N+1},u_{N+1},\tau^{*})$, with $\tau^{*}$ in (5), has marginal validity guarantees, as shown by the theorem below.
Theorem 2.1 (Marginal validity of nested conformal prediction (Angelopoulos et al.,, 2020)).
Assume an exchangeable sequence $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N+1}$ and let $\hat{Y}(\bm{x},u,\tau)$ be a set-valued predictor that satisfies (4). Furthermore, assume that $\exists\tau\in\mathbb{R}:\hat{Y}(\bm{x},u,\tau)=\mathcal{Y}$. Then, for $\tau^{*}$ in (5) and any $\alpha\in(0,1)$, the following marginal coverage guarantee holds:
@@
In the literature, several conformal prediction methods can be found that satisfy (4). For example, Romano et al., (2020) propose the following set-valued predictor, called adaptive prediction sets (APS):
@@
with $\hat{\rho}(y;\bm{x})=\sum_{y^{\prime}\in\mathcal{Y}}\hat{P}(y^{\prime}\,|\,\bm{x})\mathbbm{1}_{\hat{P}(y^{\prime}\,|\,\bm{x})>\hat{P}(y\,|\,\bm{x})}$ the probability mass of the labels more likely than $y$. At the heart of their method is the randomization term $u\cdot\hat{P}(y\,|\,\bm{x})$, which is used to achieve exact nominal coverage. In addition, by means of the cumulative distribution $\hat{\rho}(y;\bm{x})$, this method allows to adapt effectively to complex data distributions, achieving a better conditional coverage compared to alternative approaches that rely on the mode of the distribution, such as the least ambiguous classifier proposed by Sadinle et al., (2019). Angelopoulos et al., (2020) propose a variant of the above method, called the regularized adaptive prediction sets (RAPS) method. By introducing a regularization term in (6), small set sizes are encouraged, especially when facing noisy probability estimates for classes with low probability.
Alternative conformal prediction methods have been introduced, focusing on minimizing set size (Sadinle et al.,, 2019; Gao et al.,, 2025) and achieving approximate coverage across the entire feature space (Foygel Barber et al.,, 2021; Cauchois et al.,, 2021; Romano et al.,, 2019). Goren et al., (2024) and Angelopoulos et al., (2023) propose conformal prediction methods that control the risk with respect to bounded non-increasing loss functions. Interestingly, these methods are also applicable to the setting of hierarchical classification, e.g. through the use of the tree-distance loss, introduced by Bi and Kwok, (2015) as a way of evaluating set-valued predictions in multi-label classification. However, set-valued predictions of that kind lack meaningful and practical interpretation, compared to methods that rely on the traditional notion of coverage. Furthermore, these methods are not directly applicable to our work, as they do not allow the incorporation of representation complexity, given the properties of the loss functions above.
Gao et al., (2025) recently proposed a novel conformal predictor for regression, by implementing a dynamic programming algorithm that finds a union of $k$-intervals on the real line. When the output distribution is multi-modal, a union of $k$-intervals can be more informative and efficient compared to a single prediction interval (as obtained with most regression methods). Our work has a similar motivation, but adopted to hierarchical classification: when the classifier is uncertain between classes in different branches of the hierarchy, restricting the representation complexity to one is overly restrictive, while a higher complexity yields more efficient predictions.